OpenStudy (anonymous):

3HBr + Al(OH)3 ----> 3H2O + AlBr3 If 8.56 grams of HBr react with an excess of Al(OH)3 to form 8.91 grams of AlBr3, what is this reaction’s percent yield? Could somebody please help walk me through this?

OpenStudy (anonymous):

first find theoretical yield (mass of AlBr3 formed from 8.56g of HBr) just use the steps from the previous question but in this case moles of AlBr3=1/3 moles of HBr then find % yield using the formula % yield=experimental yield (given in question)/theoretical yield x 100

OpenStudy (anonymous):

wait can we start fresh and over from step 1?

OpenStudy (anonymous):

ok first step find Mr of limiting reagent (reagent not in excess in this case HBr) step 2 find moles of HBr step 3 moles of AlBr3=1/3 moles of HBr (from equation) step 4 find mass of AlBr3 this is the theoretical yield divide the yield given in the question by this yield and multiply by 100 to find %

OpenStudy (chmvijay):

still you didnt get it ????

OpenStudy (anonymous):

no I can figure it out when someone gives me the steps, but on my own I don't know where to start

OpenStudy (chmvijay):

ok then let me try it

OpenStudy (chmvijay):

since Al(OH)3 is in excess reactant we can say that limiting reagen will be HBr right

OpenStudy (anonymous):

ok seems only logical

OpenStudy (chmvijay):

yaaa now can you tell the molecular weight of HBr and AlBr3

OpenStudy (anonymous):

how do I find it?

OpenStudy (anonymous):

molar mass?

OpenStudy (chmvijay):

yaaa

OpenStudy (chmvijay):

no u find the molar mass

OpenStudy (rane):

lol vijay

OpenStudy (chmvijay):

why LOL

OpenStudy (anonymous):

molar mass of HBr is 80.91g molar mass of AlBr3 is 266.69g

OpenStudy (rane):

once u said yes next minute no...bt the answer is same

OpenStudy (chmvijay):

ok now 80.91*3 gram of HBr produces =266.69 g ram of AlBr3 then 8.56 gram of Hbr should produce = 266.69*8.56 /80.91 =28.21 gram of AlBr3 you get

OpenStudy (chmvijay):

28.21 gram of AlBr3 =for 100 % but u r getting 8.91 gram =8.91*100 /28.21 = 31.58 %

OpenStudy (chmvijay):

LOL @rane sometime it happens :) nobody is perfect

OpenStudy (rane):

i know no body is perfect calm down

OpenStudy (anonymous):

yes I get it, so if 8.56 grams of HBr react with an excess of Al(OH)3 to form 8.91 grams of AlBr3, this reaction’s percent yield is 31.58%

OpenStudy (anonymous):

Thank you!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

it seems easier now

OpenStudy (chmvijay):

you are welcome :)